system of equations
5x-2y=4
3x+y=9
the answer is in this form (?,?)

Answer by rfer(16322)

3x+y=9
y=-3x+9
---------------
5x-2(-3x+9)=4
5x+6x-18=4
11x=22
x=2
------------------
3*2+y=9
6+y=9
y=3
----------------
3*2+3=9